Using your target mass of CO2, calculate the mass of citric acid monohydrate, C6H8O7·H2O, you would need to produce it. The target mass is 0.50g and i need to solve it with a three step bracket.


1
Expert's answer
2021-01-19T04:40:56-0500

you can make three steps bracket like this

(CO2)3 (C4H8O) (H2O)


carbon di oxide mass is given = 0.50g

no. of moles of CO2 = 0.50/44 = 0.0113 moles

since we've 3 units , so total CO2 moles = 0.0340 moles


Mass of C4H8O unit = 0.0340x 72 = 2.45 g

Mass of H2O unit = 0.0340 x 18 = 0.612 g


Total mass = 0.50+ 2.45+ 0.612 = 3.562 grams